OpenStudy (anonymous):

-5(3m+6)=-3(4m-2). What is the answer show steps plz

OpenStudy (ybarrap):

$$ -5(3m+6)=-3(4m-2)\\ -15m-30=-12m+6 $$ Understand this so for?

OpenStudy (ybarrap):

*far

OpenStudy (ybarrap):

Does this make sense?